OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point:(8, -4) ; Slope: -5/6

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 8}}\quad ,&{\color{blue}{ -4}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies -\cfrac{5}{6} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

so its y-4=5/6 (x-8) ?